OpenStudy (anonymous):

Find each product 5v^2(8v+8)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$5v^2(8v+8)$$

OpenStudy (skullpatrol):

\[a \left( b +c \right)=ab+ac\]